OpenStudy (anonymous):

help, medal and fan Determine the zeros of the function x2 - 16x + 63 = 0. {7, 9} {-7, 9} {-9, 7} {-9, -7}

OpenStudy (anonymous):

@stacy.reed08 what it mean by zero function?

OpenStudy (anonymous):

i think its (7,9)

OpenStudy (anonymous):

(x-7)(x-9) = x^2-7x-9x+63 = x^2 - 16x + 63 yes the zeros are 7 and 9.

OpenStudy (anonymous):

the zeros are when the equations crosses the x-axis.

OpenStudy (anonymous):

so yes, 7 and 9 :)

OpenStudy (anonymous):

Its A!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Yes, option A is correct.

OpenStudy (anonymous):

thank you guys! i wish i can give everyone medals lol